A gentleman farmer maniacally works on academic projects to raise funds for his true passion, farming. He has identified seven tasks and developed three-time estimates (all in days) of the next big project on the farm, replacing the fence at the comedy pasture. Tasks, precedence requirements, and time estimates are shown in the table. What is the probability that the project is completed in 31 days?

Activity Optimistic time Most Likely time Pessimistic time Predecessor
Corner Posts 4 5 6 -
String Line 7 8 9 -
Dig Holes 10 12 14 Corner Posts
Plant Posts 11 14 17 String Line
Weld Top Bar 8 11 14 String Line
Paint 9 11 13 Dig Holes, Plants Posts
Tack Wire 9 12 16 Weld Top Bar
Answer #1

Answer: the probability that the project is completed in 31 days= 0.0544

Explanation:

Activity Optimistic time-a Expected completion time-m Pessimistic time-b Expected time= (a+4*m+ b)/6 Variance, (sigma)^2= (b-a/6)^2
A 4 5 6 5.00 0.11
B 7 8 9 8.00 0.11
C 10 12 14 12.00 0.44
D 11 14 17 14.00 1.00
E 8 11 14 11.00 1.00
F 9 11 13 11.00 0.44
G 9 12 16 12.17 1.36

For critical path

Paths Duration
ACF 28.00
BDF 33.00
BEG 31.17

BDF is the longest path- critical path, with duration 33 days

For probability

step 1 we will find the variance of the tasks which lie on critical path 1.56
step 2 mean project time (u) of critical path is= 33.00
step 3 Required completion time is 31
step 4 standard deviation= sqrt(variance)=   1.247
step 5 because Z= (given completion time- u)/standard deviation -1.604
step 6 P(z)= using NORM.S.DIST(z,true) 0.0544

The probability is 0.0544

Please ask, if you have any doubts through the comment section. Do rate the answer